# LetsGo: Large-Scale Garage Modeling and Rendering via LiDAR-Assisted Gaussian Primitives

### v1

@@ 4.1 LiDAR-Assisted Gaussian Splatting
Furthermore, our method also supports residual Gaussian, a hybrid representation composed of meshes and Gaussian. In scenarios of large-scale underground garages, most parts consist of continuous planes like the ground, which can be effectively utilized by a rough mesh. Meanwhile, details such as texture and glasses on the car can be effectively represented by the Gaussian. As shown in Fig. 4, by applying this hybrid representation, the memory consumption for training and model storage can be significantly reduced.
@@ 4.2 Training Details
Figure 6: Qualitative comparison between our LiDAR-GS, 3DGS, SuGaR, Mip-Splatting, NGP and F2-NeRF on the various datasets. Here, we use a superscription to denote their Gaussian primitives are initialized according to our LiDAR-assisted 3D point cloud.
Table 2: Quantitative comparison between our LiDAR-GS, 3DGS, SuGaR, Mip-Splatting, NGP and F2-NeRF on the various datasets.
@@
Ours(LiDAR-GS) 24.63 0.835 0.400 27.54 0.903 0.149 21.26 0.722 0.269
We train our model using the PyTorch Framework on NVIDIA RTX A6000 GPUs. Considering the large scale of our scene and the use of LiDAR point clouds for initialization, we adjust the scaling learning rate to 0.0015 and set the initial position learning rate for residual Gaussians primitives to 0.000016. Additionally, we increase the opacity reset interval to 2,000,000 and delay the Gaussian densification step start to 75,000. The total iteration count is empirically set to twenty times the number of captured images. We set the spherical harmonics (SH) degree to 2 and the weights for regularization terms $\lambda_{\mathrm{depth}}$ and $\lambda^{*}_{\mathrm{depth}}$ to 0.8, respectively. We also apply partition-based training to train our data in parallel for acceleration.
Block Partitioning.
As illustrated in Fig. 5, we partition each level of garages into sub-blocks with overlapped quadrilateral, typically a rectangle or a trapezoid with small internal angles. Each block covers approximately 100 square meters. This enables scene representation training at a small scale.
Edge Expansion and Point Cloud Partitioning.
Due to poor reconstruction quality at edges by Gaussian splatting based methods, we expand each quadrilateral outward by 30% to enhance the reconstruction quality. Since the collected garage point clouds are vertically aligned with the Z-axis, we only need the quadrilateral vertices for point cloud segmentation and downsample the point cloud at 4cm intervals.
Image Selection.
After processing the original fisheye images, we retain camera views with origins inside the partition. For cameras outside the partition, we project the complete and quadrilateral-partitioned meshes to the views, comparing the number of triangles visible from each view. Views with a ratio greater than a certain threshold, determined as 0.8 from experiments, are retained.
Results Merging.
After partition training, we use the original quadrilateral coordinates to segment and merge the results, achieving a complete scene reconstruction.
Figure 7: Illumination of different Gaussian splatting octree levels. The root level contains the lowest density subsample of the original Gaussian splats, and with each level, the resolution is doubled.
4.3 Experimental Results
@@
In this section, we compare our approach with the recent state-of-the-art on various challenging large-scale datasets, including our GarageWorld, KITTI-360 [33], a large outdoor street scene, and ScanNet++ [78], an indoor scene characterized by complex geometry and variable lighting. We adopt 10% images of each scene for test sets, others are training sets.
@@
We compare our method against a suite of 3DGS approaches, including 3DGS*[25], Mip-Splatting*[80], and SuGaR*[20]. They all utilized our LiDAR-derived point clouds as initialization, enabling their application to the challenging garage scenes. Additionally, our approach is benchmarked against recent implicit representation methods, namely F2-NeRF[70] and Instant-NGP[43],
Evaluation Metrics.
For quantitative comparisons, we adopt Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS)[83] as metrics to evaluate the rendering quality.
Results.
@@
As shown in Fig. 6, 3DGS* and SuGaR* tend to generate floating Gaussian points around scene surfaces, leading to blurriness and poor rendering quality. F2-NeRF performs well on the relatively small-scale scene of ScanNet++ but suffers from artifacts in the large-scale KITTI-360 dataset and blurry in our GarageWorld dataset. As for Instant-NGP, which utilizes hash encoding, its rendering results in large-scale scenes appear notably blurry. In contrast, our method shows the highest rendering quality in various scenarios. The comparison between Ours (LiDAR-GS) and 3DGS* demonstrates that our depth regularizer successfully suppresses the float Gaussian primitives, resulting in improved rendering quality. The quantitative comparisons between the state-of-the-art methods on the three datasets are presented in Tab. 2. Our method consistently achieves the best result.
5 Web-based Lightweight Renderer
Lightweight rendering is crucial for enabling real-time visualization on commonly used devices like laptops and mobile phones. It reduces the dependency on high-performance computing devices and ensures a positive user experience across various application scenarios. In the context of large-scale scenes, exemplified by the garages discussed in this paper, the 3D Gaussian splatting necessitates a substantial number of Gaussians to depict the entire scene comprehensively. However, mainstream lightweight devices currently grapple with the challenge of simultaneously loading all Gaussian splats into VRAM for rendering. To mitigate this, this paper partitioned the large garage scene into small sub-blocks. Despite this optimization, the required number of Gaussians is still significant, making real-time rendering on lightweight devices challenging.
@@ 5 Web-based Lightweight Renderer
In response to the complexities posed by large-scale scenes, recent approaches [57, 36] have turned to a Level of Detail (LOD) structure for storing and loading point clouds or textured mesh data into memory. Drawing inspiration from this LOD structure, this paper introduces a new renderer tailored to our LiDAR-assisted Gaussian primitives. We present a LOD structure for 3D Gaussian storage and rendering at different resolutions. This innovation enables real-time, high-quality rendering on lightweight devices, addressing the unique demands of large-scale scenes.
5.1 Level of Detail Structure for 3D Gaussians
Similar to Potree [57], we store the unstructured 3D Gaussians of our representation in the partitioned sub-block of the large-scale garage in a layered manner in an octree, as shown in Fig. 7. Each level of the Octree contains a subsample of the whole data at a specific resolution, with the density/resolution of the subsample decreasing with the level number. The max level of the Octree contains the original residual 3D Gaussian data at a maximum level of detail, while the lowest level denotes the root node and stores the coarsest subsample of the whole data at the lowest level of detail. Following previous works, we apply chunking, merging, and downsampling operations to create the octrees at each level.
The chunking step splits the original 3D Gaussians into cubic chunks with an appropriate chunk size, which is small enough that parallel processing for multiple chunks can be guaranteed while being large enough to contain adequate numbers of 3D Gassians to avoid massive tiny chunks. In our scenario, we divide the space of the partitioned sub-block into a chunk grid of ${128^{3}}$ and count at most 10k 3D Gaussians within each grid. Unlike previous works for point cloud LOD rendering that only need to consider the position of the 3D point, we take into account both the position and radius of the 3D Gaussians during the chunking step. A 3D Gaussian is assigned to a chunk when its center position is within this chunk and also at least 50% of its radius extends within the boundaries of the chunk, ensuring that most of the Gaussian volume is contained within the designated spatial segment. We merge adjacent chunks containing fewer than 10k 3D Gaussians to avoid sparsity within any chunk. To create the LOD structure, we downsample the leaf nodes of a higher-level octree to generate the subsample of 3D Gaussians at the lower level. This process is repeated recursively until the coarsest level.
We store the chunking information at each level, including the whole region of the garage scene, the number of 3D Gaussians, and the starting and ending address of each chunk in binary files. During the following rendering processes, these metadata are loaded into memory, assisting in deciding which parts of the 3D Gaussians to access.
5.2 Coarse-to-Fine Rendering
We implement a coarse-to-fine loading and rendering scheme for the residual 3D Gaussians to enable real-time rendering on lightweight devices. We divide the LOD structure for 3D Gaussians into two parts. Levels under four are preloaded at the beginning and kept in memory. These are primarily used for rendering the overall scene at a lower resolution. Since they store a relatively smaller number of 3D Gaussians and are frequently accessed data, keeping them continuously in memory does not significantly impact resource usage while helping avoid I/O issues associated with frequent loading. Levels higher than four are dynamically loaded into memory to compensate for scene details of the coarse levels based on the given camera parameters. To ensure the rapid rendering of the dynamically loaded data, we employ a multi-threaded asynchronous reading approach, loading each level of 3D Gaussians in parallel. Additionally, to guarantee the loading time of the new 3D Gaussian nodes at the updated viewing positions, we set an upper limit on the size of data loaded at once. This limit is set to 2 million 3D Gaussians per load in our implementation.
The above-illustrated loading and rendering parameters could be adjusted according to the memory size and computational power of different devices, thereby accommodating the rendering tasks of garage scenes on various devices. We test the performance of our renderer on different devices. On a high-performance desktop equipped with an i9-10900X CPU and a Samsung SSD T5 Disk, it takes 1.36 seconds to load approximately 2 million 3D Gaussians. On a MacBook laptop equipped with an Apple M2 CPU and APPLE SSD AP0512Z Disk, it takes 1.29 seconds to load the same 2 million 3D Gaussians. Combining these extra 3D Gaussians with the preloaded 3 million Gaussians of the scene, both devices maintain a rendering frame rate of 60 FPS.
For our hybrid mesh and residual Gaussian representation, we leverage an open-source library, Nexus [48], to store and render mesh from different resolutions, and the residual Gaussian primitives are rendered using the above illustrated LOD renderer.
6 GarageWorld and Applications
@@
Fig.  11 visualizes the GarageWorld dataset, consisting of five large-scale garages that exhibit a diverse array of types: underground garages, outdoor parking, an indoor garage with multi-floors, and an indoor garage featured with mechanical parking systems. These environments are characterized by intricate geometric structures, such as sloped surfaces with distorted lines, internal circular paths, spiral ramps, and parked vehicles. The point clouds illustrated in Fig.  11 capture the comprehensive architectures of the garages. Our neural rendering pipeline is employed to produce images from multiple perspectives, exemplifying the photorealistic quality of our rendering technique. Our neural rendering pipeline and the GarageWorld dataset are designed to support an extensive range of applications, details of which are delineated in the forthcoming sections.
@@
6.1 Data Generation and Testbed for Autonomous Driving
Our reconstructed 3D structure for large-scale garage scenes and high-quality real-time rendering assists various autonomous driving algorithms. For example, autonomous vehicles are equipped with multiple sensors, including LiDAR, IMU and cameras, to perceive the environment. The reconstructed 3D structure aids in sensor fusion algorithms by providing a diverse and accurate dataset for training. Our GarageWorld dataset includes a wide range of driving scenarios, such as tight parking spaces, multi-level structures, complex spiral and circular paths, diverse geometric layouts. This provides various solution generation for different autonomous driving applications, such as the trajectory generation for autonomous parking, path planning for navigation, etc. With the generated data, autonomous driving algorithms can be trained to handle these complex scenarios, improving their ability to operate effectively in various garage environments. Furthermore, benefiting from our reconstructed garage world by LiDAR-assisted Gaussian primitives, algorithms can be rigorously tested and validated in a controlled virtual environment before being deployed in real-world scenarios. This accelerates the development and deployment process while ensuring a higher level of safety. Additionally, the challenging garage environments, as discussed previously, with large textureless regions, repetitive patterns, transparent and reflective surfaces, etc., offer a realistic testbed for SfM and visual SLAM techniques.
6.2 Real-time Localization and Navigation
Apart from generating various training data in complex scenarios, our LiDAR-assisted garage modeling and rendering also benefit autonomous vehicle localization and navigation by providing accurate 3D references and complementing real-time sensor readings.
In the challenging indoor or underground garage environment, there is often limited or no access to GPS signals, making traditional GPS-based localization challenging. The confined spaces and structured obstacles can lead to interference for LiDAR and radar sensors, and multipath reflections in the confined spaces can distort sensor readings, making localization by LiDAR and radar less reliable. Furthermore, the low lighting in underground and indoor garages reduces camera visibility, posing challenges for purely vision-based localization. Benefiting from our Polar scanner and LiDAR-assisted Gaussian primitives, we provide accurate 3D references with colors that assist devices in accurately recognizing and matching environmental features. This integration significantly enhances localization and navigation accuracy and reliability.
Fig.  9 presents an example of real-time localization and navigation. When the task is triggered, our localization system correlates an image captured at the current location with our reconstructed colored 3D models and computes a relative pose. Then, our navigation system generates an optimal path connecting the current location and destination in the complex garage environment. During the driving process, our 3D model and real-time rendering complement the vehicle camera readings and broaden the field of view of the vehicle, providing a comprehensive understanding of the environment, thus enabling safe driving. Our web-based rendering engine empowers our system to represent expansive 3D garage maps and navigation trajectories on lightweight devices with limited computational capacity, thus delivering smooth real-time interactions and superior rendering efficacy.
6.3 Autonomous Vehicle Parking
Our diverse and complex garage environment provides various challenging parking scenarios with different parking trajectories for vehicle parking algorithm training. During deployment, based on the vivid rendering output of the garage, autonomous driving systems analyze available parking spaces, generate a trajectory, and guide the vehicle into these spots along the trajectory.
Apart from this trajectory planning, our LetsGo pipeline for garage modeling and rendering can also assist autonomous vehicles in safe parking by complementing sensor readings. As discussed previously, due to the low light conditions in garage environments, real-time sensor observations by vehicle-mounted cameras may not observe the surrounding scene accurately. The vehicle-equipped LiDAR or Radar sensors may suffer inaccuracies due to reflections in the confined space. Our detailed 3D model helps identify and classify obstacles within the garage, ensuring safe and efficient parking. As shown in Fig. 8, our method generates rear view images with a wide field of view, encapsulating drivable area and obstacles (e.g., walls). This allows accurate distance measurement between the vehicle and the obstacle, avoiding collisions. The proposed lightweight real-time rendering allows the vehicle to dynamically assess its surroundings and make quick decisions, increasing safety and navigating the vehicle through tight spaces.
@@ 6.3 Autonomous Vehicle Parking
Figure 10: VFX demonstration. Through an analysis of the animation in the reference video, we manually extract the poses of several keyframes, which enable our system’s renderer to generate corresponding video segments. Our 3D garage modeling and rendering also enables motion blur rendering, producing realistic visual effects.
6.4 VFX Production
Our reconstructed garage datasets and real-time rendering can significantly contribute to VFX productions. For example, iconic scenes like the highway chase in ”The Matrix Reloaded” are captivating and impressive. Filming such scenes can involve shooting or simulating extreme locations using visual effects. Our LetsGo pipeline can achieve high-quality modeling and rendering of complex garage scenes. Moreover, real-time rendering enables the manipulation of various factors, such as exposure time and the significance of motion blur, which can be challenging to adjust in actual filming scenarios. Our GarageWorld serves as a valuable foundation for creating lifelike backgrounds or integrating CGI (Computer-Generated Imagery) elements seamlessly into live-action footage.
To demonstrate the capabilities of our approach, we create a VFX video production using our reconstructed garage models (Fig. 10) and the supplementary video. In this demonstration, we download a video depicting highway chase scenes with dynamic viewing angles of a rapidly moving motorcycle. By manually extracting the camera trajectory, we align it with one of our garage scenes and render corresponding images. We also synthesize motion blur during the rendering process, creating compelling visual effects of a chase inside a garage.
Moreover, our 3D garage model serves as a versatile background that can be overlaid with special effects, such as shots of cars, for movie scene creation. With the advantage of real-time rendering, VFX artists can explore various camera angles and compositions before finalizing the shot. This eliminates the need for extensive reshooting, saving valuable time and resources in the production pipeline. Additionally, the appearance consistency between the real and virtual components can be guaranteed, enhancing the overall believability of the scenes.
Figure 11: A gallery showcasing the rendering results of our pipeline across various scenes in GarageWorld. Our LiDAR-assisted Gaussian primitives enable photorealistic rendering of expansive garages. The enlarged images illustrate the fidelity of our high-resolution point clouds and captured imagery.

### v2

@@ 4.3 Web-based Lightweight Renderer
Converting Multi-Resolution Gaussians into Octree Structure
Our multi-resolution Gaussian representation is well-suited for storage within an Octree structure, with Gaussians at each level efficiently stored in the nodes at the corresponding depth of the Octree. We begin by constructing an Octree of depth $L$, where the root node host level 0 and the leaf nodes contain the finest level, $L-1$. Intermediate levels are stored at the inner nodes of the Octree. We outline the conversion of our Gaussian representation into the Octree format in Algorithm 1.
Coarse-to-Fine Rendering
Rendering large-scale garage scenes requires managing a substantial number of Gaussians, often exceeding in-memory storage capacity. To address this, we implement a coarse-to-fine loading and rendering scheme that facilitates real-time rendering on resource-constrained devices. Initially, low-resolution Gaussian levels are preloaded into memory at viewer initialization. These levels occupy minimal storage and are accessed frequently during subsequent rendering cycles. As shown in Fig. 4, to render high-frequency details, our system dynamically selects high-resolution octree chunks from disk based on the viewer’s frustum and the conditions set by our LOD rendering strategy, as outlined in Equation 8. The rendering process begins with a traversal of the entire octree, determining each node’s visibility based on the center position of the node’s bounding box. A second traversal then loads visible Gaussian data into a buffer, continuing until either the traversal is complete or the number of loaded Gaussians reaches the maximum capacity. The buffer is updated by replacing preloaded Gaussians with newly loaded high-resolution Gaussians, which are then dispatched to a WebGL worker for rendering. Recent works, such as OCT-GS [61] and LoG [70], can also be structured into an octree format. However, OCT-GS fails to represent the entire scene at certain levels, making it unsuitable for our lightweight rendering strategy. Additionally, LoG does not provide a solution for lightweight rendering.
5 Experimental Results
@@ 5.1 Training Details
Figure 7: Qualitative comparison between our LOD-LiDAR-GS, 3DGS, Mip-Splatting, F2-NeRF, NGP and Splatfacto on the various datasets. Here, we use a superscription to denote their Gaussian primitives are initialized according to our LiDAR-assisted 3D point cloud.
Table 2: Quantitative comparison among our LOD-LiDAR-GS, 3DGS, Mip-Splatting, F2-NeRF, NGP, Splatfacto, OCT-GS, and LoG-GS on the various datasets.
@@
Splatfacto* 25.45 0.793 0.261 28.35 0.906 0.065 14.76 0.481 0.381
OCT-GS* 25.19 0.782 0.315 31.55 0.941 0.063 21.52 0.692 0.322
LoG-GS* 21.53 0.737 0.278 27.80 0.907 0.086 18.62 0.671 0.300
Ours(LOD-LiDAR-GS) 25.77 0.812 0.210 29.19 0.927 0.064 24.53 0.811 0.167
Ours(Web) 22.56 0.730 0.211 22.79 0.824 0.140 19.20 0.586 0.247
All experiments are conducted on an NVIDIA RTX A6000 GPU using the PyTorch Framework. Given the large scale of our scenes and the initialization with LiDAR point clouds, we set the scaling learning rate to 0.0015 and the initial position learning rate for Gaussian primitives to 0.000016. We disable the opacity reset option and delay the start of the Gaussian densification step to 75,000 iterations. The total number of iterations is empirically set to twenty times the number of captured images. We use a spherical harmonics (SH) degree of 2 and set $\lambda_{\mathrm{depth}}$ to 0.8.
Block Partitioning and Merging.
To accelerate our model training through parallel execution, we employ a partition-based strategy. As illustrated in Fig. 6, each garage floor is divided into sub-blocks using overlapping quadrilaterals, typically rectangles or trapezoids with small internal angles. Each block covers approximately 100 square meters, enabling efficient small-scale training for scene representation. To address the poor reconstruction quality at the edges, common in Gaussian splatting-based methods, we expand each quadrilateral outward by 30%. This expansion enhances the overall reconstruction quality. After training each partition, we merge the results using the original quadrilateral coordinates, achieving a complete scene reconstruction. This method ensures high-quality rendering while optimizing the training process.
Training Image Selection and Pre-processing.
Regarding training view selection for each block, camera views originating within the partition bounds are retained. Furthermore, for camera views positioned outside the partition, we project the mesh in the partition block and the original mesh onto each camera view and compute the projected pixel overlap ratio between the quadrilateral-partitioned mesh and the whole mesh. Views with a ratio exceeding a threshold of $0.8$ are retained. The original Gaussian framework does not support direct input of fisheye images. To retain more information from the images, we avoided traditional fisheye distortion correction. Instead, we split a single fisheye image into five pinhole images. Specifically, the fisheye image is projected onto a hemispherical surface. Then, five virtual cameras, each oriented at 45 degrees to the top, bottom, left, right, and straight ahead, reproject the hemispherical image into their respective views. This process generates new images and extrinsic parameters. Additionally, since the downward-facing view includes part of the capturing device, we use a mask to obscure this portion of the image.
5.2 Comparison
@@
We compare our approach with state-of-the-art methods and concurrent works on various challenging datasets, including our GarageWorld, KITTI-360 [44] (large outdoor street scene), and ScanNet++ [92] (indoor scene characterized by complex geometry and variable lighting). For each dataset, we use 10% of the images as test sets and the remaining 90% as training sets.
@@
We compare our method against a suite of 3DGS approaches, including 3DGS* [34], Mip-Splatting* [94], and Splactfacto* [76], as well as recent concurrent work, including OCT-GS* [61] and LoG* [70]. Additionally, we benchmark our approach against recent implicit representation methods, namely F2-NeRF [81] and Instant-NGP [53]. To extend recent 3DGS-based methods to large-scale garage scenes, we initialize them with our LiDAR-derived point clouds instead of SFM point clouds, enabling their application to challenging garage environments. Methods utilizing our LiDAR-derived point clouds are denoted with an asterisk (*).
Figure 8: Qualitative comparison of our method and concurrent works, including recent OCT-GS* and LoG* methods, on the GarageWorld and UAVD4L datasets.
Qauantitative Comparison Results.
For quantitative comparisons, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS)[96] as metrics to evaluate rendering quality. The results of these comparisons are presented in Tab. 2. Our LOD-LiDAR-GS method consistently achieves the best results on large-scale scene datasets, such as our GarageWorld dataset and the KITTI-360 dataset. On small-scale scene dataset ScanNet++, our LOD-LiDAR-GS approach achieves the second-best results. From the last two rows of Tab. 2, our web renderer demonstrate comparable performance on lightweight devices (MacBook Air with M2 clip).
Figure 9: Qualitative ablation study on depth regularization within our large-scale GarageWorld dataset clearly demonstrates improvements in resolving the floater problem on the ground.
Figure 10: Ablation study on our LOD rendering strategy demonstrates that our real-time LOD renderer operates approximately 3-4 times faster than the original 3DGS SIBR viewer when tested on an NVIDIA RTX 3090 GPU.
Qualitative Comparison Results.
@@
As shown in Fig. 7, both 3DGS* and Splactfacto* tend to generate floating Gaussian points around scene surfaces, resulting in blurriness and poor rendering quality. F2-NeRF performs well on the relatively small-scale ScanNet++ dataset but suffers from artifacts in the large-scale KITTI-360 dataset and blurriness in our GarageWorld dataset. Instant-NGP, which utilizes hash encoding, produces notably blurry results in large-scale scenes. In contrast, our method demonstrates the highest rendering quality across various scenarios. We also compared our method with concurrent works such as OCT-GS* and LoG*. As shown in Fig. 8, our approach achieves the best rendering quality on both our GarageWorld and the UAVD4L datasets (recently proposed and used in LoG*). Our web-based results also show comparable visual quality.
Rendering Performance of Web-based Renderer.
Additionally, we evaluate the performance of our web-based renderer on various devices. On a high-performance desktop equipped with an i9-10900X CPU and a Samsung SSD T5 Disk, it takes 1.36 seconds to load approximately 2 million 3D Gaussians. On a MacBook laptop with an Apple M2 CPU and Apple SSD AP0512Z Disk, the load time for the same number of 3D Gaussians is 1.29 seconds. Combining these newly loaded 3D Gaussians with an additional 3 million pre-loaded Gaussians, both devices maintain a rendering frame rate of 60 FPS.
5.3 Ablation Study
In this section, we perform ablation studies to validate the effectiveness of each component of our approach.
Depth Regularization.
Using only photometric loss constraints can easily produce artifacts, such as floaters, in large-scale scene reconstruction. To address this, we introduce an additional depth regularization term to enforce geometric constraints. As shown in Fig. 9, we evaluated the impact of our depth regularization module across various scenes. Our results demonstrate that the LOD-LiDAR-GS method effectively mitigates floaters on the ground when depth regularization is applied.
LOD Rendering Strategy.
Next, we ablate the LOD rendering strategy. Based on our multi-resolution representation, we select Gaussian subsets from various resolution levels according to depth ranges, following the LOD rendering strategy outlined in Equation 8. This strategy significantly reduces the complexity of sorting, projection, and accumulation operations involved with 3D Gaussians, thereby accelerating the rendering speed. As shown in Fig. 10, our real-time viewer, integrated with the LOD rendering technique, is about 3-4 times faster than 3DGS’s SIBR viewer with similar rendering quality.
Random-Resolution-Level (RRL) Training Strategy.
During training, directly applying the LOD rendering strategy to render images and optimize the loss can lead to overfitting at certain resolution levels in specific scene regions. We propose a random-resolution-level (RRL) training strategy to address this issue. As shown in Fig. 11, our RRL training strategy mitigates the overfitting problem and remains robust to variations in the distribution of input cameras.
Figure 11: Qualitative ablation study on the random-resolution-level (RRL) training strategy reveals that our RRL approach effectively addresses the overfitting issue at a certain level, influenced by the camera distribution.
6 GarageWorld
@@
We introduce the first large-scale garage dataset, GarageWorld, captured with our Polar device. This dataset comprises eight large garages, featuring six underground garages, one outdoor parking lot, and one multi-floor indoor garage. These environments present diverse and intricate geometric structures, including sloped surfaces with distorted lines, internal circular paths, spiral ramps, and parked vehicles, as detailed in Table 1. Figure 5 visualizes the 3D models and rendered images of a subsection of these garages using our method. Our rendering pipeline and the GarageWorld dataset are designed to support a wide range of applications, which are further detailed in the following sections.
Data Generation and Testbed for Autonomous Driving.
Our developed 3D reconstruction of large-scale garage scenes with enhanced real-time rendering can support various autonomous driving algorithms. The GarageWorld dataset provides diverse driving scenarios, including tight parking spaces and complex geometric layouts, which are instrumental for training algorithms in autonomous parking and navigation path planning. Equipped with sensors such as LiDAR, IMU, and cameras, autonomous vehicles utilize these 3D structures for effective sensor fusion, enhancing their operational efficiency in varied garage environments. Additionally, our LetsGo pipeline further aids vehicle parking by supplementing sensor data, particularly in low-light garage conditions where camera and radar readings may be compromised. Our LiDAR-assisted Gaussian primitives allow rigorous testing in a virtual setting, prompting safe and efficient parking solutions. Fig. 12 illustrates how our approach generates rear-view images, allowing vehicles to accurately estimate distances and navigate tight spaces, thereby avoiding collisions and accelerating the development and deployment of safe autonomous driving technologies.
Real-time Localization and Navigation.
Our LiDAR-assisted garage modeling and rendering significantly enhance autonomous vehicle localization and navigation by providing accurate 3D references that complement real-time sensor readings. In challenging indoor or underground garages, where GPS signals are typically unavailable, traditional localization methods encounter difficulties. Confined spaces and structured obstacles can cause interference for LiDAR and radar sensors, and multipath reflections can distort sensor readings. Low lighting conditions also reduce camera visibility, complicating vision-based localization. By integrating our Polar scanner with LiDAR-assisted Gaussian primitives, we offer accurately colored 3D references that enhance feature recognition and matching, significantly improving localization and navigation reliability. Figure 13 presents an example of real-time localization and navigation. Our system correlates captured images with our colored 3D models to compute relative poses and generate optimal paths through complex environments. During driving, our approach augments camera views and extends the vehicle’s field of view, ensuring comprehensive environmental awareness and safer navigation. Additionally, our web-based rendering engine represents expansive 3D garage maps and navigation trajectories on lightweight devices with limited computational capacity, delivering smooth real-time interactions and superior rendering performance.
@@
VFX Production.
Our reconstructed garage datasets and real-time rendering capabilities significantly enhance VFX production by providing a robust foundation for creating realistic backgrounds and seamlessly integrating CGI elements into live-action footage. Inspired by high-adrenaline sequences like the highway chase in ”The Matrix Reloaded,” our LetsGo pipeline excels in modeling and rendering complex garage scenes. This capability allows for precise control over cinematic elements such as exposure time and motion blur, which are often challenging to manage on set. We created a VFX video production to showcase our technology using our reconstructed garage models (Fig. 14) in the supplementary video. We downloaded a video depicting highway chase scenes with dynamic viewing angles of a rapidly moving motorcycle. By extracting and aligning camera trajectories with our models and synthesizing motion blur, we created immersive visual effects that convincingly depict a chase within a garage environment. Our approach supports the exploration of diverse camera angles and compositions in real-time while ensuring visual consistency between real and virtual components, significantly reducing the need for reshoots and streamlining the production process. This integration empowers VFX artists to enhance the believability of scenes while conserving time and resources.
@@
Figure 14: VFX demonstration. Through an analysis of the animation in the reference video, we extract the poses of several keyframes, which enable our system’s renderer to generate corresponding video segments. Our 3D garage modeling and rendering also enables motion blur rendering, producing realistic visual effects.